Min Z = 6X1 + 4x2 Subject to Xi + 2x2 > 2 -X1 + 2x2 5 4 3x1 + 2x2 < 12 X1, X2 > 0
Answer #1

Ano Min z = 6 xxt 442 Subotom 8172307 32 - 3 24+242 3. 4 321+272512 - 0 . X1, X₂ 7,0 Solved coing Graphically 0 21+2.42 = 2 -Values of Function At (012) Z = 6 27+ 4X2 =16 XOlt (4x2) = 8 1238 At could 2 =621+4 X2 =66x0) 44x1) = 4 2-4 At 1-1,3/2) Za 6x